Maltose + H2O⟷2 Glucose

Calculate the equilibrium constant under standard transformed conditions (
1
Expert's answer
2020-08-09T08:29:53-0400

Maltose + H2O⟷2 Glucose


Gibbs Free Energy Change under standard transformed conditions = -15.5 kj /mol.


So , K = Equillibrium Constant = 10-15.5 *1000 / 2.303 * 298 * 8.314.


So , K = 10-2.71 = 2 * 10-3 . Answer .